I need help in these equation, thank you
you are varnishing the background for a rectangular mural. The base of the mural is 7/2 meters and the height of the mural is 5 meters. How many cans of varnish will you need if each can covers 10 square meters?
A)4 cans of varnish
B) 7 cans of varnish
C) 18 cans of varnish
D) 2 cans of varnish

Answer by stanbon(75887)

I need help in these equation, thank you
you are varnishing the background for a rectangular mural. The base of the mural is 7/2 meters and the height of the mural is 5 meters. How many cans of varnish will you need if each can covers 10 square meters?
-------
Area of the mural = (7/2)*5 = 35/2 = 17.5 sq. meters
--------
# of cans = 17.5/10 = 1.75 cans
-----
Cheers,
Stan H.
===========================
A)4 cans of varnish
B) 7 cans of varnish
C) 18 cans of varnish
D) 2 cans of varnish